# Mastering Board Games by External and Internal Planning with Language Models

### v1

@@ G.1 Chess
51…Qh3 A long maneuvering phase of the game results in a four queens endgame! MAV–MCTS is completely winning because of a very weak black king. 52. Qbb8 Qe6+ 53. b3 f5 54. gxf6 Bxf6 55. Qb7+ Be7 56. Qbxe7+ Qxe7 57. Qxe7+ Kh8 58. Qf6+ Kg8 59. Qxf4 exf4 60. a5 g5 61. a6 Kg7 62. a7 Kf7 63. a8=Q g4 64. Qd5+ Ke7 65. Qd6+ Kf7 66. Qf6+ Ke8 67. Nc5 gxf3 68. Qg7 f2 69. Qd7+ Kf8 70. Ne6+ Kg8 71. Qg7# 1 – 0
@@
19. Ngxe6 MAV–IS sacrifices a knight to keep black king in the center. Nxe6 20. Bxb5+ Bc6 21. Nxe6 Bxb5 22. Re1 Bd6 23. Nxg7+ Kd7 24. Rxe5 Bxe5 25. Qxe5 MAV–IS correctly estimated that the vulnerable black king and dark squares are worth the sacrificed exchange. Qc5
@@ G.2 Chess960
Game 1 Setting the board on fire Date: 2024-11-14
@@
MAV–MCTS cannot prevent black pawn from promoting on move 10! But… 11. Ba4 a1=Q 12. Bxe8 Bxf2+ 13. Kxf2 Qa5
@@
Game 2 Accelerated Benko gambit Date: 2024-11-14
@@
Game 3 Delayed Benko gambit Date: 2024-11-14
@@
Game 4 Not fearing ghosts Date: 2024-11-14
@@
Game 5 Finishing games in style Date: 2024-11-14
@@
Game 6 A knight on the rim… Date: 2024-11-14
@@
Game 7 Complications never end Date: 2024-11-14
@@
Game 8 Only active pieces count Date: 2024-11-14
@@
22. c4 MAV–MCTS point is finally revealed – the black rook is isolated from the rest of the board! Qg5 23. Qxg5 Nxg5 24. Kc2 Nf3 25. Kc3 a5 26. Rxb7 Nxh2 27. Rf2 Kxb7 28. Rxh2 b5 29. a3
@@
The black rook gets trapped by a pawn. 29…Rxd3+ 30. Kxd3 Rxb3+ 31. Kd4 bxc4 32. Rxh7 Kc6 33. Ne4 Rd3+ 34. Kxc4 Rxa3 35. Nc3 Ra1 36. Rxf7 Kd6 37. Rf6+ Ke5 38. Rxg6 Kf5 39. Rg8 Ke5 40. Rg4 Ke6 41. Rh4 a4 42. Rh2 Rg1 43. Ne2 Rb1 44. Rh5 a3 45. Ra5 Ra1 46. Nd4+ Kd6 47. Rd5+ Ke7 48. Kb3 a2 49. Nc6+ Ke6 50. Rh5 Rg1 51. Kxa2 Rg2+ 52. Ka3 Kf6 53. Rh1 Rg3+ 54. Ka4 Kf5 55. Ka5 Ra3+ 56. Kb6 Ra8 57. Rh5+ Ke4 58. Rh6 Ra1 59. Rh8 Rb1+ 60. Ka6 Rb2 61. Na7 Ra2+ 62. Kb7 Rb2+ 63. Kc8 Rb1 64. Rh6 Kf4 65. Ra6 Rh1 66. Kd8 Ke3 67. Ra5 Rh7 68. Ra1 Kd3 69. Nc8 Kc3 70. Rg1 Rh5 71. Kd7 Rh7+ 72. Ke6 Kc2 73. Rg8 Kd2 74. Rg5 Rh6+ 75. Kd7 Rh7+ 76. Kd6 Rh4 77. Rg8 Ke3 78. Na7 Kf4 79. Kc6 Rh6+ 80. Kb5 Rh1 81. Kc6 Rh6+ 82. Kb5 Ke4 83. Kb4 Kd5 84. Nc8 Rh1 85. Rg5+ Kd4 86. Kb5 Rh8 87. Na7 Ke3 88. Kb6 Rh1 89. Rg8 Ra1 90. Rg4 Rc1 91. Rg6 Rb1+ 92. Kc6 Ra1 93. Rg7 Ke2 94. Rg2+ Kf1 95. Rg7 Rb1 96. Rg4 Rc1+ 97. Kb7 Rb1+ 98. Ka6 Ra1+ 99. Kb5 Rxa7 100. Rg6 Ra8 101. Kc5 Kf2 102. Rb6 Ke3 103. Rb1 Ra5+ 104. Kd6 Ke4 105. Rb4+ Kf5 106. Rb6 Ra1 107. Kd5 Rd1+ 108. Kc4 Ra1 109. Kd4 Ra4+ 110. Kd5 Ra5+ 111. Kd4 Ra1 112. Rh6 Ra4+ 113. Kd3 Kg5 114. Rh7 Ra1 115. Ke4 Kg6 116. Rh2 Re1+ 117. Kd3 Ra1 118. Ke4 Kf7 119. Rh7+ Kg6 120. Rh2 (threefold repetition). An extraordinary game despite the outcome! 1/2 – 1/2
Game 9 Returning knight Date: 2024-11-14
@@
Game 10 Harrys’ deadlock Date: 2024-11-14
@@
Game 11 Fianchettoing the rook Date: 2024-11-14
@@
Game 12 Bishops vs knights Date: 2024-11-14
@@
Bishop pair against two knights imbalance guarantees a complex middlegame 24. Kh2 Rg8 25. e4 g5 26. fxg6 Rxg6 27. Qh3 Kd8 28. Bd1 Rg7 29. Qe3 Qa5 30. a3 Rg8 31. Re1
@@
Game 13 The last resource Date: 2024-11-14
@@
Appendix H Opening Book

### v3

@@ F.1 Chess
51…Qh3 A long maneuvering phase of the game results in a four queens endgame! MAV–MCTS is completely winning because of a very weak Black’s king. 52. Qbb8 Qe6+ 53. b3 f5 54. gxf6 Bxf6 55. Qb7+ Be7 56. Qbxe7+ Qxe7 57. Qxe7+ Kh8 58. Qf6+ Kg8 59. Qxf4 exf4 60. a5 g5 61. a6 Kg7 62. a7 Kf7 63. a8=Q g4 64. Qd5+ Ke7 65. Qd6+ Kf7 66. Qf6+ Ke8 67. Nc5 gxf3 68. Qg7 f2 69. Qd7+ Kf8 70. Ne6+ Kg8 71. Qg7# 1 – 0
@@
19. Ngxe6 MAV–IS sacrifices a knight to keep Black’s king in the center. Nxe6 20. Bxb5+ Bc6 21. Nxe6 Bxb5 22. Re1 Bd6 23. Nxg7+ Kd7 24. Rxe5 Bxe5 25. Qxe5 MAV–IS correctly estimated that the vulnerable Black’s king and dark squares are worth the sacrificed exchange. Qc5
@@ F.2 Chess960
Game 24 Setting the board on fire Date: 2024-11-14
@@
MAV–MCTS cannot prevent Black’s pawn from promoting on move 10! But… 11. Ba4 a1=Q 12. Bxe8 Bxf2+ 13. Kxf2 Qa5
@@
Game 25 Accelerated Benko gambit Date: 2024-11-14
@@
Game 26 Delayed Benko gambit Date: 2024-11-14
@@
Game 27 Not fearing ghosts Date: 2024-11-14
@@
Game 28 Finishing games in style Date: 2024-11-14
@@
Game 29 A knight on the rim… Date: 2024-11-14
@@
57. Rd5 The Black’s knight is trapped on the edge of the board! f5 58. Rxf5 g4 59. Rg5 Rg8 60. Rxg8+ Kxg8 61. Kd2 Kf8 62. Ke3 Kg7 63. Kf4 Nc2 64. Kxg4 Kf6 65. Be8 Ne3+ 66. Kf4 Nd5+ 67. Ke4 Nc7 68. Bg6 Na6 69. Kd4 Kg5 70. Kc4 Kh6 71. Kb5 Nc7+ 72. Kc6 Na6 73. Kb5 Nc7+ 74. Kc5 Kg7 75. Bd3 Kh6 76. Kd6 Ne8+ 77. Ke7 Ng7 78. Bg6 Nxh5 79. Bxh5 Kg5 80. Bd1 Kh4 81. Ke6 Kg5 82. Ke5 Kg6 83. b4 Kg5 84. b5 Kh6 85. Kf6 Kh7 86. b6 Kg8 87. b7 Kh7 88. b8=R Kh6 89. Rh8# 1 – 0
Game 30 Complications never end Date: 2024-11-14
@@
Black’s queen is nearly trapped. 18…Qd6 19. Ne2 Nc5 20. Ra3 MAV–MCTS finds the time for a prophylactic move in the middle of the complications. e4 21. fxe4 fxe4 22. a6 Bb6 23. axb7 Bxa7 24. c4 Rxb7
@@
Game 31 Only active pieces count Date: 2024-11-14
@@
22. c4 MAV–MCTS point is finally revealed – the Black’s rook is isolated from the rest of the board! Qg5 23. Qxg5 Nxg5 24. Kc2 Nf3 25. Kc3 a5 26. Rxb7 Nxh2 27. Rf2 Kxb7 28. Rxh2 b5 29. a3
@@
The Black’s rook gets trapped by a pawn. 29…Rxd3+ 30. Kxd3 Rxb3+ 31. Kd4 bxc4 32. Rxh7 Kc6 33. Ne4 Rd3+ 34. Kxc4 Rxa3 35. Nc3 Ra1 36. Rxf7 Kd6 37. Rf6+ Ke5 38. Rxg6 Kf5 39. Rg8 Ke5 40. Rg4 Ke6 41. Rh4 a4 42. Rh2 Rg1 43. Ne2 Rb1 44. Rh5 a3 45. Ra5 Ra1 46. Nd4+ Kd6 47. Rd5+ Ke7 48. Kb3 a2 49. Nc6+ Ke6 50. Rh5 Rg1 51. Kxa2 Rg2+ 52. Ka3 Kf6 53. Rh1 Rg3+ 54. Ka4 Kf5 55. Ka5 Ra3+ 56. Kb6 Ra8 57. Rh5+ Ke4 58. Rh6 Ra1 59. Rh8 Rb1+ 60. Ka6 Rb2 61. Na7 Ra2+ 62. Kb7 Rb2+ 63. Kc8 Rb1 64. Rh6 Kf4 65. Ra6 Rh1 66. Kd8 Ke3 67. Ra5 Rh7 68. Ra1 Kd3 69. Nc8 Kc3 70. Rg1 Rh5 71. Kd7 Rh7+ 72. Ke6 Kc2 73. Rg8 Kd2 74. Rg5 Rh6+ 75. Kd7 Rh7+ 76. Kd6 Rh4 77. Rg8 Ke3 78. Na7 Kf4 79. Kc6 Rh6+ 80. Kb5 Rh1 81. Kc6 Rh6+ 82. Kb5 Ke4 83. Kb4 Kd5 84. Nc8 Rh1 85. Rg5+ Kd4 86. Kb5 Rh8 87. Na7 Ke3 88. Kb6 Rh1 89. Rg8 Ra1 90. Rg4 Rc1 91. Rg6 Rb1+ 92. Kc6 Ra1 93. Rg7 Ke2 94. Rg2+ Kf1 95. Rg7 Rb1 96. Rg4 Rc1+ 97. Kb7 Rb1+ 98. Ka6 Ra1+ 99. Kb5 Rxa7 100. Rg6 Ra8 101. Kc5 Kf2 102. Rb6 Ke3 103. Rb1 Ra5+ 104. Kd6 Ke4 105. Rb4+ Kf5 106. Rb6 Ra1 107. Kd5 Rd1+ 108. Kc4 Ra1 109. Kd4 Ra4+ 110. Kd5 Ra5+ 111. Kd4 Ra1 112. Rh6 Ra4+ 113. Kd3 Kg5 114. Rh7 Ra1 115. Ke4 Kg6 116. Rh2 Re1+ 117. Kd3 Ra1 118. Ke4 Kf7 119. Rh7+ Kg6 120. Rh2 (threefold repetition). An extraordinary game despite the outcome! 1/2 – 1/2
Game 32 Returning knight Date: 2024-11-14
@@
Game 33 Harrys’ deadlock Date: 2024-11-14
@@
Game 34 Fianchettoing the rook Date: 2024-11-14
@@
Game 35 Bishops vs knights Date: 2024-11-14
@@
Bishop pair against two knights imbalance guarantees a complex middlegame. 24. Kh2 Rg8 25. e4 g5 26. fxg6 Rxg6 27. Qh3 Kd8 28. Bd1 Rg7 29. Qe3 Qa5 30. a3 Rg8 31. Re1
@@
Game 36 The last resource Date: 2024-11-14
@@
Appendix G Opening Book